When a LR circuit with L=3mH is connected in series to an AC source of rms voltage 30V a maximum RMS current of 6A is observed at frequency (500/π)Hz. If LR circuit is now connected to a battery of emf 21V And internal resistance of 3ohm the current will be
Ans 3A
1
Expert's answer
2020-12-02T09:46:02-0500








Let the RMS current be "I" "I^2 = \frac{V^2}{R^2+(wL)^2}"

"w=2(pi)f"



"I=6A"

"f=\frac{500}{pi}"

Where "pi=\frac{22}{7}"

Solving the above equations we get R= 4 ohm.

When the LR circuit is connected to battery of emf 21V and internal resistance 3 ohm then new current is i .

"i =\frac{21}{4+3}"

"i= 3A"